OpenStudy (anonymous):

2. Set up the following word problem as a proportion and solve. Then write your answer as a complete sentence. A health department estimates that 8 vials of malaria serum will treat 100 people. At this rate, how many vials are required to treat 175 people?

OpenStudy (anonymous):

We know that 8 vials are good for 100 people, so to know what a single vial will treat we need to divide 100 by 8. 100/8 = 12.5 (one vial will treat 12.5 people) We now divide 175 (total people needing treatment) by 12.5 (amount of people treated by 1 vial) and that gives us : 175/12/5 = 14